If I have a 80ml solution that is 20% acid by volume, how many mL should I remove and replace with pure 100% acid to obtain a final solution of 80ml that is 40% acid by volume? What is the equation?
Answer by stanbon(75887)

If I have a 80ml solution that is 20% acid by volume, how many mL should I remove and replace with pure 100% acid to obtain a final solution of 80ml that is 40% acid by volume? What is the equation?
-----
Equation:
acid - acid + acid = acid
0.20*80 - 0.20x + x = 0.40*80
---
Multiply thru by 100 to get:
20*80 - 20x + 100x = 40*80
80x = 20*80
x = 20 ml (amt. to remove and replace)
================
Cheers,
Stan H.